40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 231 of 354  Not logged in ELOG logo
ID Date Author Type Category Subject
  6242   Wed Feb 1 17:00:57 2012 steveUpdateIOOlaser is back ON

Quote:

 

The 2W PSL laser is turned off.  The danger laser lights are not illuminated at the entry doors because of malfunctioning electronic circuit!!!

Laser safety glasses are still required!  Other lasers are in operation!

 BEN fixed the interlock.  The laser is turned ON. Thanks for all, Rich and Sam who came over to help. Atm1

All emergency shut- off switches, lights and door indicators are working at this moment. More about this tomorrow.

Atm2, PSL enclosure interlock jungle without REAL schematic drawing.....at this point.... We all agreed it is easier to redo the hole thing than find the problem

Atm3, Emergency shut off switches and illuminated signs from  entry doors to AC on-off box  ( Use this switches in emergency ONLY,  otherwise leave alone , even it is labeled obsolete !)

Summery: I still do not really know what was wrong.

 

Attachment 1: P1080525.JPG
P1080525.JPG
Attachment 2: P1080514.JPG
P1080514.JPG
Attachment 3: P1080518.JPG
P1080518.JPG
  6241   Tue Jan 31 17:13:49 2012 steveUpdateIOOonly the PSL laser is off

 

The 2W PSL laser is turned off.  The danger laser lights are not illuminated at the entry doors because of malfunctioning electronic circuit!!!

Laser safety glasses are still required!  Other lasers are in operation!

  6240   Tue Jan 31 14:58:30 2012 kiwamuUpdateIOOlaser shuts down

[Steve/ Kiwamu]

 We found that the laser had completely shut off for ~ 4 hours even with all the PSL doors closed.

We are guessing it is related to the interlock system and Steve is working on it to fix it.

Quote from #6239

 The 2W Innilight shutdown shut when I opened side door for safety scan. This was not a repeatable by opening -closing side doors later on. Turned laser on, locked PMC and MC locked instantly. The MC was not locked this moring and it seemed that the MC2 spot was still some high order mode

like yesterday. MC lock was lost when the janitor bumped something around the MC.

 

  6239   Tue Jan 31 08:44:10 2012 steveUpdateIOOlaser shuts down

 

 The 2W Innilight shutdown shut when I opened side door for safety scan. This was not a repeatable by opening -closing side doors later on. Turned laser on, locked PMC and MC locked instantly. The MC was not locked this moring and it seemed that the MC2 spot was still some high order mode

like yesterday. MC lock was lost when the janitor bumped something around the MC.

Attachment 1: laseroff.png
laseroff.png
  6238   Mon Jan 30 23:10:02 2012 kiwamuUpdateSUSdrift mon script needs to be fixed

[Rana / Kiwamu]

We tried to set some parameters for the suspension drift monitor but the old matlab script, which automatically sets the values, didn't run because it uses the old mDV protocol.

The attached link below is a description about the script.

https://wiki-40m.ligo.caltech.edu/Computers_and_Scripts/All_Scripts#Drift

It needs to be fixed or upgraded by pynds.

  6237   Mon Jan 30 16:18:51 2012 steveUpdatePEMRoscolux colored film transmittance at 1064 nm

 

 Roscolux filter films  #74 night blue,  0.003" thick  and  #26 light red, 0.002" thick  were measured in the beam path of  ~6 mm diameter,  1W 1064 nm .

T 90%  + - 5% at 0-30 degrees of  incident angles and R ~10 % 

These sandwitched thin films of policarbonate-polyester filters are not available in thicker forms. Rosco is recommending them to be cooled by air if used in high power beam.

These filters did not get warm at all in 1W, so absorption must be very small.

  6236   Mon Jan 30 08:17:06 2012 steveUpdateIOOPMC

Quote:

I have realigned the beam pointing to PMC. The transmitted light increased from 0.74 to 0.83.

The misalignment was mainly in pitch.

 The PMC pointing has changed, so MC is resonating in high order modes.

Attachment 1: sickPMC.png
sickPMC.png
  6235   Fri Jan 27 17:16:05 2012 kiwamuUpdateLSChypothetical glitch scenario

Here is a hypothetical scenario which could make the glitches in the LSC error signals. It can be considered as a 4 step phenomenon.

        (1) up conversion noise due to a large motion at 3 Hz

   => (2) rms level exceeds the line width (a.k.a. linear range) in some LSC sensors

   => (3) unlocks some of the DOFs  in a moment

   => (4) glitches due to the short unlock.

- - plan  - -

 In order to check this hypothesis the low finesse PRMI must serve as a good test configuration.

What I will do is to gradually decrease the offset in MICH such that the finesse of PRMI becomes higher.

And at each different finesse I will check the spectra, glitch rate, and etc.

Quote from #6231

        Low finesse PRMI      

In this configuration NO glitches ( a high speed signal with an amplitude of more than 4 or 5 sigma) were found when it was locked.

Is it because I didn't use AS55 ?? or because the finesse is low ??

Also, as we have already known, the up conversion noise (#6212) showed up -- the level of the high frequency noise are sensitive to the 3 Hz motion.

  6234   Fri Jan 27 16:55:28 2012 JenneUpdateGreen LockingY-green realigned

The Yarm green laser really wanted to lock on a 01/10 mode, so Kiwamu suggested I go inside and realign the green beam to the arm.  I did so, and now it's much happier locked on 00 (the Yarm is resonating both green and IR right now).

  6233   Fri Jan 27 13:13:03 2012 JenneUpdateSUSITMs tripped

Sitting down to start cavity measurements, I found both ITMs tripped.  It must have happened a while ago (I didn't bother to check dataviewer trends) because both had rms levels of <5 counts, so they've had a while to sit and quiet down.

  6232   Fri Jan 27 09:05:41 2012 steveUpdateSUSOSEM locking plunger

Quote:

Our existing 300 series SS plungers from McMastercar #8476A43 are silver plated as Atm2 shows.

Problems:  1, they become magnetized after years being close to the magnets

                     2, they oxidize by time so it is hard to turn them

                    

I looked around to replace them.

Titanium body, nose and beryllium copper spring. None magnetic for UHV enviorment.

Can be made in 7 weeks at an UNREASONABLE $169.00 ea at quantity of 50

 In order to get a better price from Vlier's Tom Chen I changed Ti body back to SS304L-siver plated and music wire spring. The price is still ~$120 ea. at quantity 50

I will talk to Mike G about modifying the  McMaster plunger with a hex nut.

  6231   Fri Jan 27 06:07:47 2012 kiwamuUpdateLSCglitch hunting

I went through various IFO configurations to see if there are glitches or not.

Here is a summary table of the glitch investigation tonight. Some of the cells in the table are still not yet checked and they are just left blank.

IFO configuration  Yarm
Xarm
MICH
Half PRMI
low finesse PRMI
PRMI (carrier)
PRMI (sideband)
DRMI
AS55 NO NO NO   up conversion noise glitch glitch glitch
REFL11 NO NO NO   up conversion noise
glitch glitch glitch
REFL33 NO NO NO   - glitch glitch glitch
REFL55 NO NO NO   up conversion noise
glitch glitch glitch
REFL165 NO NO NO   - glitch glitch glitch
POX11 - NO NO     glitch glitch glitch
POY11 NO - NO     glitch glitch glitch
POP55 - -            
                 

 

        Low finesse PRMI      

The low finesse PRMI configuration is a power-recycled MIchelson with an intentional offset in MICH to let some of the cavity power go through MICH to the dark port.

To lock this configuration I used ASDC plus an offset for MICH and REFL33 for PRCL.

The MICH offset was chosen so that the ASDC power becomes the half of the maximum.

In this configuration NO glitches ( a high speed signal with an amplitude of more than 4 or 5 sigma) were found when it was locked.

Is it because I didn't use AS55 ?? or because the finesse is low ??

Also, as we have already known, the up conversion noise (#6212) showed up -- the level of the high frequency noise are sensitive to the 3 Hz motion.

  6230   Fri Jan 27 05:21:43 2012 kiwamuUpdateGreen Lockingfine alignment of the Y end green setup

I did a fine alignment on the Y end green setup. The green light became able to be locked again.

Quote from #6227

 The alignment is finished after the realization that the 3rd steering mirror had to be adjusted too.

  6229   Thu Jan 26 19:28:02 2012 kiwamuUpdateIOOPMC low transmission

After I recovered the lock of PMC, I found that the PMC transmission was quite low. It was about 0.26 in the EPICS display.

I zeroed the PSL temperature feedback value which had been -2.3 and then the PMC transmission went back to a normal value of 0.83.

I believe it was because the PSL was running with two different oscillation modes due to the big temperature offset.

  6228   Thu Jan 26 15:35:23 2012 JenneUpdateIOOMC ~1Hz badness

The mode cleaner is super unhappy.  It's rocking around at ~1Hz.

I turned off the WFS and turned them back on after the MC was locked, and it seems a little happier now.  At least it's not falling out of lock ~1/minute.

  6227   Thu Jan 26 10:17:01 2012 steveUpdateGreen Lockinggeen pointing into y arm is realigned

Quote:

I  placed an other Y2-LW-1-2050-UV-45P/AR steering mirror into the beam path of the green beam launching in order to avoid the ~30 degrees use of the 45 degrees mirror. The job is not finished.

 The alignment is finished after the realization that the 3rd steering mirror had to be adjusted too.

The input power increased from 1.2 to 1.4 mW

Attachment 1: P1080513.JPG
P1080513.JPG
  6226   Thu Jan 26 08:36:38 2012 steveUpdateSAFETYevacuation drill

It started with fire alarm test yesterday at 14:50 All alarms are  functioning VERY loud and their flashers are bright. Evacuation drill followed. We assembled at north west corner of the 40m building and counted 6 heads.

Nobody was left sleeping inside. Bob carried the success report of the drill to PMA office immediately.

  6225   Thu Jan 26 06:09:52 2012 kiwamuUpdateGreen Lockingnoisy AS55

During the Y arm ALS I found that the noise of the AS55 demod signal was worse than that of POY11 in terms of the Y arm displacement.

There is a bump from 500 mHz to 100 Hz in the AS55 signal while POY11 didn't show such a structure in the spectrum.

 

The plot below is the noise spectra of the Y arm ALS. The arm length was stabilized by using the green beat-note fedback to ETMY.

In this measurement, POY11 and AS55 were served as out-of-loop sensors, and they were supposed to show the same noise spectra.

In the plot It is obvious that the AS55 curve is louder than the POY curve.

Yarm_ALS_2012Jan25.png

  6224   Thu Jan 26 05:40:10 2012 kiwamuUpdateLSCglitch in the analog demodulated signals

Indeed the glitches show up in the analog demodulated signals. So it is not an issue of the digital processing.

With an oscilloscope I looked at the I/Q monitor outputs of the LSC demodulators, including REFL11, REFL33, REFL55, POY11, AS55 while keep locking the carrier-resonant PRMI.

I saw some glitches in REFL11, REFL55 and AS55. But I didn't see any obvious glitches in REFL33 and PO11 because the SNR of those signals weren't good enough.

 


(some example glitches)

The attached plot below is an example shot of the actual signals when the carrier resonant PRMI was locked.

The first upper row is the spectrogram of REFL11_I, REFL55_I, REFL33_I and AS55_Q in linear-linear scale.

The second row shows the actual time series of those data in unit of counts.

The bottom row is for some DC signals, including REFLDC, ASDC and POYDC.

 

glitch.png

You can see that there are so many glitches in the actual time series of the demod signals (actually I picked up the worst time chunk).

It seems that  most of the glitches in REFL11, REFL33 and AS55  coincide.

The typical time scale of the glitches was about 20 msec or so.

Note that the PRMI was locked by REFL33 and AS55 as usual.

  6223   Wed Jan 25 17:32:03 2012 steveUpdateGreen Lockinggeen pointing into y arm is misaligned

I  placed an other Y2-LW-1-2050-UV-45P/AR steering mirror into the beam path of the green beam launching in order to avoid the ~30 degrees use of the 45 degrees mirror. The job is not finished.

  6222   Wed Jan 25 15:04:41 2012 JenneBureaucracySAFETY40m SOS supplies moved off of cleanroom flow bench

Bob, Callum and Daphen noted that our keeping a JDSU HeNe (max power <4mW) is against somebody's SOP.  So I cleared everything that relates to 40m SOS suspending to the bottom shelf of the 2nd cabinet in the cleanroom (the back set of cabinets nearest the flow benches).  The door has a nifty label.  Things that are in there include:

HeNe

HeNe mount

QPD and micrometer mount

microscope and micrometer mount

iris

Al beam block

Magnet gluing fixture

dumbbell gluing fixture

The electronics that we use (HeNe's power supply, 'scope, QPD readout) are still on the roll-y thing under the flow bench.

  6221   Wed Jan 25 02:59:46 2012 kiwamuUpdateGreen LockingY arm ALS noise budget

Surprisingly increasing the gain of the whitening filter didn't improve the noise curve.

It suggests that the ADC noise is not the limiting factor below 10 Hz.

Quote from #6220

Though, it turned out that the MFD's ADC is now limiting the noise in a frequency band of 200 mHz - 5 Hz.

So tonight I will increase the gain of the whitening filter to push down the ADC noise more.

 

  6220   Tue Jan 24 18:11:13 2012 kiwamuUpdateGreen LockingY arm ALS noise budget

I did some more stuff for the Y arm ALS and updated the noise budget:

After the works, the rms displacement improved a little bit, so it is now at 24 pm in rms.

Though, it turned out that the MFD's ADC is now limiting the noise in a frequency band of 200 mHz - 5 Hz.

So tonight I will increase the gain of the whitening filter to push down the ADC noise more.

 

Yarm_ALS_2012Jan19.png

 

(What I did)

 + added the DAC noise and comparator noise based on measurements.

 + redesigned the servo filter shape to suppress the seismic noise below 10 Hz.

 The attached plot below shows the newly designed open loop transfer function together with the old one for a comparison.

UGF is at 120 Hz and the phase margin is about 27 deg.

  newservo.png

  • FM7 = resonant gain (17)
  • FM6 = resonant gain (3)
  • FM5 = zero(1) * pole(500)
  • FM4 = pole(1) * zero(40.) * 40.
  • FM3 = pole(1) * zero(40.) * 40.
  • FM2 = pole(0.001)*zero(1)*1000.
  6219   Tue Jan 24 13:36:05 2012 ZachBureaucracyGeneralIf I'm Peter Pan...

jamie_rufio.png

JA - MIE - RO!

  6218   Mon Jan 23 23:12:00 2012 kiwamuUpdateIOOPMC realignment

I have realigned the beam pointing to PMC. The transmitted light increased from 0.74 to 0.83.

The misalignment was mainly in pitch.

  6217   Mon Jan 23 15:43:47 2012 JenneUpdateIOOPMC realignment

Quote:

I realigned the incident beam to PMC at 23:30. The transmitted light went up from 0.78 to 0.83.

 Do we have PSL pos and ang QPD trends?  We should start watching them, because the PMC drifted back down to 0.76 transmission, ~3.5 days after Kiwamu realigned it (his elog is from last Thurs).  Not so awesome.

I walked through the control room just now and found both PMC and MC unlocked.  They're both locked now, but with PMC transmission 0.76, MC transmission ~24,500.

  6216   Fri Jan 20 17:05:59 2012 ranaUpdateGreen LockingY arm ALS : time series

                        One of my goals this week is to get people to make plots with physical units:

That ALS plot would be 5x cooler if the POY11 signal could be in meters instead of counts or cubits.

  6215   Fri Jan 20 16:24:50 2012 kiwamuUpdateGreen LockingY arm ALS : time series

Here is a new time series plot showing how stably ALS can control the arm length.

In the middle of the plot the cavity length was held at the resonance point for ~ 2 min. and then it passed through the resonance point to show the full shape of the PDH signal.

Apparently the PDH signal is now quieter than before (#6133)

time_series.png

Quote from #6214

One of my goals in this week is : measurement of the current best ALS noise budget.

 

  6214   Fri Jan 20 15:59:02 2012 kiwamuUpdateGreen LockingY arm ALS noige budget

One of my goals in this week is : measurement of the current best ALS noise budget.

Last night I took a new noise spectra of the Y arm ALS, which is shown in the attached figure below.

The displacement of the arm cavity observed from the IR PDH is at 66 pm in rms. In the measurement the arm length was stabilized with the ALS technique.

 

Yarm_ALS_2012Jan19.png

 

  6213   Thu Jan 19 23:34:52 2012 kiwamuUpdateIOOPMC realignment and HEPA

I realigned the incident beam to PMC at 23:30. The transmitted light went up from 0.78 to 0.83.

Also I decreased the HEPA level down to 20 % for the night time locking.

  6212   Wed Jan 18 16:31:10 2012 kiwamuUpdateLSCestimation of optical length between PRM and scattering object

I searched for a scattering body in the REFL path.

According to the result the REFL path on the AS table is innocent.

 

The idea of the search method is given as follows:

  •   Put a 1/10 ND attenuator at the origin of the REFL path on the AS table.
  •   Of course this reduces the signal level by the same factor of 10 in the REFL11_I demod signal.
  •   If the scattering body is in the REFL path the up conversion noise will be smaller by a factor of 100 because the scattered light go across the attenuator twice.

 

 The attached plot below is the spectra of REFL11 with the 1/10 attenuator at the origin of the REFL path when the beam is single-bounced from PRM.
In the measurement PRM_POS was driven at 10 Hz with an amplitude of 1700 all the time. This is exactly the same situation as that explained in the previous elog entry (#6211).
You can see that the up conversion noise level also decreased by the same factor of 10, which suggests there are no scattering object in the REFL path.
Note that the data with the attenuator in place is intentionally scaled by multiplying a factor of 10 for comparison.
ND1atten.png
 

Quote from #6211

Assuming that PRM is interfering with some other optics, I have estimated the optical distance between PRM and an object that interferes with PRM.

The optical distance is estimated to be 9.5 +/- 0.5 m.

If we believe this number the object is most likely outside of the vacuum chambers.

 

  6211   Wed Jan 18 14:28:36 2012 kiwamuUpdateLSCestimation of optical length between PRM and scattering object

Assuming that PRM is interfering with some other optics, I have estimated the optical distance between PRM and an object that interferes with PRM.

The optical distance is estimated to be 9.5 +/- 0.5 m.

If we believe this number the object is most likely outside of the vacuum chambers.

 

 (The measurement)

  In order to estimate the optical length between PRM and a scattering body, I swept the frequency of the main laser by actuating on the MC length.
With the sweep, the laser frequency go across some fringes and basically it allows us to estimate the FSR of a very low finesse cavity formed by PRM and the scattering body.
Therefore we get the the optical distance based on the resultant FSR.
 
  The measurement goes as follows:
  1.  Preparation : calibration of the MC2 actuator as a frequency actuator (for more details, see the next section)
  2.  Set the interferometer to the single-bounce configuration such that the beam directly is reflected back from PRM
  3.  Take spectra of REFL11_I without driving any optics. This spectra tells us how quiet the noise normally is.
  4.  Drive MC2_POS at 10 Hz with an amplitude of 10000 counts so that we can see the high frequency up conversion noise
    • The frequency was chosen such that the excitation is out of the local damping bands
    • The amplitude was chosen to be as big as possible until the MC unlocked
    • With this drive, the laser frequency should change by 20 MHz peak-peak at 10 Hz.
  5.  Record the noisy spectrum when the MC2_POS was driven.
  6.  Drive PRM instead of MC2 at 10 Hz.
    • Adjust the amplitude of the excitation such that the cut-off frequency of the up conversion noise matches with that of the MC2 driven case.
    • The amplitude was found to be 1700 - 2000 counts, this uncertainty is currently limiting the precision of the optical distance estimation.
    • With this amount of the drive, PRM moves by 0.8 um peak-peak at 10Hz.
  7. Estimate the optical length based on the amount of the drives for PRM and MC2.
    • Estimate the FSR using the following relation df/FSR = dx/ (lambda/2). => FSR = 17 MHz
    • Since FSR = c/ (2L),  L = c/(2 FSR) = 9.5 m or so
scattering.png

 

 

(Calibration of the MC2 actuator)

 To do the measurement described above, the MC2 actuator must be calibrated in terms of a frequency actuator.
I did the same old technique (#4721): lock a cavity, adjust the UGF as low as possible, and shake an actuator of interest.
This time I used the half-PRM (PRM + ITMY) for this measurement.
The actuator responses are calibrated from that of displacement to frequency by using df/f = dx/L and assumed that L = 6.760 (#4585).
Also the PRM actuator was measured such that we can use this as a reference since we already know the response in displacement (#5637).
 The attached plot below is the actual responses that I measured yesterday. The y-axis is calibrated to Hz/counts.
 

 

calibration.png

 

Quote from #6202

Is PRM making some fringes with some other optics ??

  6210   Wed Jan 18 12:38:44 2012 steveUpdatePEMAcrylic plexiglass transmittance

Acrylite-Gp.pptx

Transparent- clear plexyglass from tree different sources were measured in 1064 and 532 nm light.

Samples: a, clear Acrylic-GP 0F00  from Ridout Plastics in thickness 0.7" ,  made by  Evonic Ind

                   b, clear cast acrylic from Mc Master Carr in thickness 0.94" , likely  made by Reynolds-Cast

                   c, clear cell cast  plexyglass from Delvie's Plastics - Utah in thickness  0.93" , maker not known

PMC reflected beam was used at 92 mW and 6 mm diameter at incident angle 0-25 degrees.

All tree samples agreed on Transmittance of ~90%, Reflectivity ~3-4% and calculated Adsorption ~6-7%

 

Transparent Colored Acrylic orange-amber #2422   from www.eplastics.com in 0.12" thickness gave  T 96%,  R 1% and  Ab-calc ~3% in the beam of 92 mW 1064  nm at 6 mm diameter.

 

Transparent , colored   Light Red #26 thin film filter   policarbonate-polyester   0.002" thick   from Roscolux measured T 81% of 115 mW 1064 nm

 

Now I changed power meter FieldMate to Ophir and the light source to laser pointer 2.2 mW  ~532 nm  with 1-2 mm beam diameter.

Orange - amber #2422  sample, 0.12" thick,  T 1% ,  R 4%  and  Ab-calculated ~95%, estimated visibility  ~50% It does cut out the green at this low power level.

 Light red #26  sample  T 0.5%  at 2.5 mW of 532 nm . The transparent green is not visible.  The softening point of this sandwiched polycarbonate-polyecter filter is 160C. Estimated VLT of this film ~40%

 

SUMMERY:

Clear and colored acrylics'  @ 1064 nm  transmittance 90% or higher  regardless of thickness. Softenig point 115 degrees C

Colored acrylic and colored policarbonate film are adsorbing the low power green and they  transmit the 1064nm beam.

Options to consider: a, acrylic laser safety shield liner of  0.125" thick inside of 1" thick clear acrylic  box, OD +5 @1064 and OD +4 @ 532nm,  amber color VLT 27%,  150$/sqft

                                      b, thin metal liner for 1" wall acrylic box, VLT 0%

                                     

 

Attachment 2: roscogel_red#26_film.pdf
roscogel_red#26_film.pdf
  6209   Wed Jan 18 12:36:26 2012 kiwamuUpdateLSCwiped a steering mirror on the REFL path

I wiped both surfaces of the REFL second steering mirror.

However no improvements. The glitches still remain.

 

(Pic.1 before wiping, Pic.2 after wiping)

DSC_3861_small.jpgDSC_3863_small.jpg

Quote from #6206

 Last night I found that there were many dust particles on the second steering mirror in the REFL path on the AS table.

  6208   Tue Jan 17 19:07:47 2012 ranaUpdateLSCglitch hunting in REFL RFPDs : strange

 Another possibility is that there is some beam clipping of the REFL beam before it gets to the PD. Then there could be a partial reflection from that creating a spurious interference. Then it would only show the fringe wrapping if you excite the scatterer or the PRM.

  6207   Tue Jan 17 16:09:20 2012 kiwamuUpdateCDSawg not working on the c1sus machine

Actually awg works fine without any problems when the excitation channels belong to the c1lsc machine.

It seems that the awg doesn't inject signals on the channels of the c1sus machine, for example C1:SUS-BS_LSC_EXC and so on.

Quote from #6204

AWG is not working. This needs to be fixed.

  6206   Tue Jan 17 13:47:40 2012 kiwamuUpdateLSCdirty steering mirror in the REFL path

 Last night I found that there were many dust particles on the second steering mirror in the REFL path on the AS table.

Looking at it through an IR viewer, I saw the REFL beam hitting one of the biggest dust particles on that mirror.

This dust particle maybe causing the glitches or maybe not.

Anyway because it's always better to have clean mirrors, I will wipe the steering mirror in this evening and check the presence of the glitches again.

Quote from #6202

The REFL11 and REFL55 demod signals show high frequency noise depending on how big signals go to the POS actuator of PRM.

Is PRM making some fringes with some other optics ??
 

  6205   Tue Jan 17 03:10:27 2012 kiwamuConfigurationIOOrotated lambda/2 plate

I have slightly rotated the lambda/2 plate, which is used for attenuating the REFL beam's power on the AS table

because the plate had been at an unusual angle for investigation of the glitches since last Thursday.

It means the laser power going to the coating thermal noise setup has also changed. Just keep it in mind.

Quote from #6198

So today we set up the Jenny RC temperature setup to lock the LWE NPRO to the RC and then set up the beat note with the IFO REFL beam on the AS table. By using the 2 laser beat, we are avoiding the VCO phase noise issue which used to limit the PSL frequency noise at ~0.01 Hz/rHz. To do this we have reworked some of the optics on the PSL and AS tables, but I think its been done without disturbing the beams for the regular locking. Beat note has been found, but the NPRO has still not been locked to the RC - next we setup the lockin amp, dither the PZT, and then use the New Focus lock box to lock it to the RC.

  6204   Tue Jan 17 02:44:59 2012 kiwamuUpdateCDSawg not working

AWG is not working. This needs to be fixed.

I could set the channel and the parameters in the AWGGUI screen, but it never inject signals to the realtime system.

  6203   Tue Jan 17 02:27:49 2012 kiwamuUpdateLSCfringe tests : all the suspensions are innocent

I did a quick test to check a hypothesis that PRM is interfering with some other optics in the single bounce configuration.

I shook all the suspensions (except the MC mirrors) at 3 Hz in POS, PIT and YAW with an amplitude of 1000 counts.

No effects were found in the REFL demod signals.

So it is NOT a fringe effect caused by the other suspended mirrors.

Quote from #6202

The REFL11 and REFL55 demod signals show high frequency noise depending on how big signals go to the POS actuator of PRM.

Is PRM making some fringes with some other optics ??   

 

  6202   Tue Jan 17 01:02:07 2012 kiwamuUpdateLSCglitch hunting in REFL RFPDs : strange

A very strange thing is going on.

The REFL11 and REFL55 demod signals show high frequency noise depending on how big signals go to the POS actuator of PRM.

This noise shows up even when the beam is single-bounced back from PRM ( the rest of the suspensions are misaligned) and it's very repeatable.

Any idea ?? Am I crazy ?? Is PRM making some fringes with some other optics ??

 

 

(background)

 The most annoying thing in the central part locking is glitches showing up in the LSC error signals (#6183).
The symptom is that when the motion in PRCL at 3 Hz becomes louder, somehow we get glitches in both the MICH and PRCL error signals.
In the frequency domain, those glitches are mostly contribute to a frequency band of about 30 - 100 Hz.
Last Thursday Koji and I locked the half PRM (PRMI with either ITMX or ITMY misaligned) to see if we still have the glitches in this simpler configuration.
Indeed there were the same kind of glitches --a loud 3 Hz motion triggers the glitches.
It was shown particularly in the REFL11 signal but not so much in the REFL33 while AS55 didn't show any glitches.
 
 
(Still glitches even in the single bounce beam)
   We were suspecting some kind of coupling from a beam jitter, so that the 3 Hz motion somehow brings the beam spot to a bad place somewhere in the REFL paths.
I misaligned all the suspensions except for PRM such that the beam directly bounces back from PRM and go to the REFL port.
Indeed there still were glitches in the REFL11 and REFL55 demod signals. It showed up once per 30 sec or so and pushes up the noise floor around 30 - 100 Hz.
There might be a little bit of glitches also in the REFL33, but the ADC noise floor and the expected glitch noise level were comparable and hence it was difficult to see the glitches in REFL33.
 

(Glitch is related to the PRM POS actuation)

 In the single-bounce configuration I started shaking the PIT and YAW motions of PRM at 3 Hz using the realtime LOCKIN oscillator to see if I can reproduce the glitches.
However no significant glitches were found in this test.
Then I started shaking the POS instead of the angular DOFs, and found that it causes the glitches.
At this point it didn't look like a glitch any more, it became more like a stationary noise.
 
 The attached screen shot is the noise spectrum of the REFL11_I.
The red curve is the one taken when I injected the 3 Hz excitation in POS by the LOCKIN oscillator.
The excitation is at 3 Hz with an amplitude of 1000 counts.
As a comparison I plotted the same spectrum when no excitation was injected and it is plotted in pink.
 

 PRMsingle_bounce.png

It seems there is a cut off frequency at 100 Hz.
This frequency depends on the amplitude of the excitation -- increasing the amplitude brings the cut off frequency higher.
This noise spectrum didn't change with and without the oplevs and local damping.
 

(Possible scenario)

A possible reason that I can think of right now is : PRM is interfering with some other optics for some reason.
But if it's true, why I didn't see any fringes in the AS demod signals in the half PRM configuration ?
 

Quote from #6183

 We tried to figure out what is causing spikes in the REFL33 signal, which is used to lock PRCL.

No useful information was obtained tonight and it is still under investigation.

  6201   Sun Jan 15 12:18:00 2012 DenUpdateAdaptive Filteringrunning time

 In order to figure out what downsampling ratio we can take, we need to determine the running time of the fxlms_filter() function. If the filter length is equal to 5000, downsampling ratio is equal to 1, number of witness channels is 1 then with ordinary compilation without speed optimization one call runs for 0.054 ms (milli seconds). The test was done on the 3 GHz Intel processor. With speed optimization flags the situation is better

-O1 0.014 ms, -O3 0.012 ms

However, Alex said that speed optimization is not supported at RCG because it produce unstable execution time for some reason. However, by default the kernel should optimize for size -Os. With this flag the running time is also 0.012 ms. We should check if the front-end machine compilers indeed use -Os flag during the compilation and also play with speed optimization flags. Flags -O3 and -Os together might also give some speed improvement.

But for now we have time value = 0.012 ms as running time for 5000 coefficient filter, 1 witness channel and downsample ratio = 1. Now, we need to check how this time is scaled if we change the parameters.

5000 cofficients - 0.012 ms

10000 coefficients - 0.024 ms

15000 coefficients - 0.036 ms

20000 coefficients - 0.048 ms

We can see that filter length scaling is linear. Now we check downsampling ratio

ratio=1 - 0.048 ms

ratio=2 - 0.024 ms

ratio=4 - 0.012 ms

Running time on the dependance of downsample ratio is also linear as well as on the dependence of the number of witness channels and degrees of freedom. 

If we want to filter 8 DOF with approximately 10 witness channels for each DOF, then 5000 length filter will make 1 cycle for ~1 ms, that is good enough to make the downsample ratio equal to 4.

Things get a little bit complicated when the code is called for the first time. Some time is taken to initialize variables and check input data. As a result the running time of the first cycle is ~0.1 ms for 1 DOF that is ~10 times more then running time of an ordinary cycle. This effect takes place at this moment when one presses reset button in the c1oaf model - the filter becomes suspended for a while. To solve this problem the initialization should be divided by several (~10) parts.

  6200   Sun Jan 15 11:40:30 2012 DenUpdateAdaptive Filteringdownsampling

Here for the downsampling process we use a low-pass Bessel digital filter of order 6, normalized cut-off frequency = 0.1. In the plot presented below we compare the results with downsampling ratio = 1, 2, 4.

2048_downsampling.png

We can see that increasing the downsampling ratio, we increase the error of the filter. Moreover, the error at some particular frequency f seems to depend on the ratio f/Fs, where Fs - sampling frequency (2048 Hz) devided by downsampling ratio. Error is the same for all curves below 1 Hz but then begins to increase as we increase the sownsampling ratio. In order to figure out what the problem is - mistake in the filter code, inaccurate upsample algorithm or this is NLMS particularity, I've changed sampling frequency in the chans/daq/C1PEM.ini and C1IOO.ini files from 2048 Hz to 512 Hz for corresponding channels. Now, we compare the error from the filter working with 2048 Hz frequency, downsampling ratio = 4, low-pass filter = Bessel of order 6, normalized cut-off = 0.1 and filter working with 512 Hz sampling frequency, without downsampling and with corresponding Bessel low-pass filter with normalized frequency 0.4.

2048_512.png

MC_F measurement at 2048 Hz was done during the day, for that reason red curved is slightly higher then green in the resonance frequencies. But still we can see that these two cases are very much alike. For this reason, it seems that NLMS filter works better with higher sampling frequencies.

  6199   Sun Jan 15 10:28:02 2012 DenUpdateAdaptive Filteringdelays

We can account for delays in the oaf system by compensating it in the adaptive path of the filter. But using only this procedure is not enough. Parameters mu and tau should be chosen accurately:

w = (1 - tau) * w;

w += mu * dw / norm;

NLMS algorithm without considering delays works well for mode cleaner length and gur1 seismometer signals, significantly reducing MC_F  with parameters mu=1, tau=0. These parameters are considered because nlms algorithm should converge with the highest speed when mu=1. However, if the system has a delay so at time moment n:

error_signal [n] = desired_signal [n] - filter_output [n-delay];

then the adaptive filter diverges for the same parameters mu=1 and tau=0 even for delay=1. For that reason we make the same calculations with tau = 1e-4 and tau = 1e-2 without reducing mu conserving the adaptation rate and get the same result as nlms algorithm without delays. Next figure shows MC_F signal, error after applying e-nlms filter with tau=1e-4 and tau=1e-2. "e-" is added to show that  a small number (epsilon) is added to the norm of the signal in order to prevent the filter from diverging in the beginning of the process when the norm is not well-determined yet.

2048_tau.png

The test was done offline with the sampling frequency 2048 Hz, without downsampling and any filters. We can see that tau=1e-4 is still not enough, tau=1e-3 or tau=1e-2 is as good as nlms without delays, tau=1e-1 and high are also bad.

Correctly choosing tau we have some freedom for delay compensation in the adaptation path. This is important as we do not know exactly what is the delay in the real system. We can measure it approximately. In order to figure out the range of reasonable delay errors we make a test with delay = 1, but to the adaptation path we give delays from 0 to 10. It turns out that adaptation path delays greater then 5 make the filter diverge, delays in the range 0-3 produce a reasonable error. In the figure below errors with adaptation path delays = 1 (correct) and 3 are presented.

2048_delays.png

  6198   Sat Jan 14 00:50:08 2012 rana, kojiConfigurationIOOTowards coating thermal noise measurement with RefCav / MC beat

Koji asked aloud tonight if we could measure the coating thermal noise of the refcav optics by beating the refcav light with the MC_TRANS light. Then we looked at our calculations for the noises:

Displacement noise of T=200ppm silica/tantala coating on a 1" silica substrate with a 300 micron beam spot = 1e-18 * sqrt(100 Hz / f) m/rHz.

Displacement noise from coating thermal in the MC is roughly smaller by the beam size ratio (1.8 mm / 0.3 mm). Some differences due to 3 mirrors and more layers on MC2 than the others, but those are small factors.

So, the frequency noise from the refcav should by larger than the MC thermal noise by a total factor of (1.8 / 0.3) * (13 m / 8 inches) ~ 400.

Another way to say it is that the effective strain noise in the RC is (1e-18 / 0.200) = 5e-18 /rHz. This translates into (5e-18 * 13) = 6.5e-17 m/rHz in the MC. (in frequency noise its 1.5 mHz/rHz).

I have measured the frequency noise in the LLO MC to be at this level back in 2009, so it seems possible to use our RC + MC to measure coating thermal noise by the length amplification factor and compete with Frank+Tara.

 

So today we set up the Jenny RC temperature setup to lock the LWE NPRO to the RC and then set up the beat note with the IFO REFL beam on the AS table. By using the 2 laser beat, we are avoiding the VCO phase noise issue which used to limit the PSL frequency noise at ~0.01 Hz/rHz. To do this we have reworked some of the optics on the PSL and AS tables, but I think its been done without disturbing the beams for the regular locking. Beat note has been found, but the NPRO has still not been locked to the RC - next we setup the lockin amp, dither the PZT, and then use the New Focus lock box to lock it to the RC.

You might think that its hard to measure this since the MC has ~1 MHz frequency fluctuations and we want to measure down to 1e-4 Hz. But, in fact, we can just use a 200 m MFD with a LT1128 preamp. Then we use the MFD to stabilize the MC length to the refcav and just use the control + error signal of the MFD setup as the coating thermal noise measurement.

 

Note: Beat found at ~40deg for the aux laser. The aux laser is on but the shutter is closed.
The AS camera seems to be hosed. Need a bit of alignment. (KA) ==> Fixed. (Jan 15)

  6197   Fri Jan 13 17:40:38 2012 ZachUpdateRF Systemfoam box and temp controller taken off of PSL table

I stole the foam EOM box and the temperature controller circuit from the PSL table so that I could continue with the RAM measurements here at the ATF.

That is all.

  6196   Fri Jan 13 16:16:05 2012 Leo SingerUpdateStewart platformFlexure type for leg joints

I had been thinking of using this flexure for the bearings for the leg joints, but I finally realized that it was not the right type of bearing.  The joints for the Stewart platform need to be free to both yaw and pitch, but this bearing actually constrains yaw (while leaving out-of-plane translation free).

  6195   Fri Jan 13 00:51:40 2012 Leo SingerUpdateStewart platformFrequency-dependent requirements for Stewart platform

Below are revised design parameters for the Stewart platform based on ground motion measurements.

The goal is that the actuators should be able to exceed ground motion by a healthy factor (say, two decades in amplitude) across a range from about .1 Hz to 500 Hz.  I would like to stitch together data from at least two seismometers, an accelerometer, and (if one is available) a microphone, but since today this week I was only able to retrieve data from one of the Guralps, I will use just that for now.

The spectra below, spanning GPS times 1010311450--1010321450, show the x, y, and z axes of one of the Guralps.  Since the Guralp's sensitivity cuts off at 50 Hz or so, I assumed that the ground velocity continues to fall as f-1, but eventually flattens at acoustic frequencies.  The black line shows a very coarse, visual, piecewise linear fit to these spectra.  The corner frequencies are at 0.1, 0.4, 10, 100, and 500 Hz.  From 0.1 to 0.4 Hz, the dependence is f-2, covering the upper edge of what I presume is the microseismic peak.  From 0.4 to 10 Hz, the fit is flat at 2x10-7 m/s/sqrt(Hz).  Then, the fit is f-1 up to 100 Hz.  Finally, the fit remains flat out to 500 Hz.

ground_velocity_spectrum.png

Outside this band of interest, I chose the velocity requirement based on practical considerations.  At high frequencies, the force requirement should go to zero, so the velocity requirement should go as f--2 or faster at high frequencies.  At low frequencies, the displacement requirement should be finite, so the velocity requirement should go as f or faster.

The figure below shows the velocity spectrum extended to DC and infinite frequency using these considerations, and the derived acceleration and displacement requirements.

requirements.png

As a starting point for the design of the platform and the selection of the actuators, let's assume a payload of ~12 kg.  Let's multiply this by 1.5 as a guess for the additional mass of the top platform itself, to make 18 kg.  For the acceleration, let's take the maximum value at any frequency for the acceleration requirement, ~6x10-5 m/s2, which occurs at 500 Hz.  From my previous investigations, I know that for the optimal Stewart platform geometry the actuator force requirement is (2+sqrt(3))/(3 sqrt(2))~0.88 of the net force requirement.  Finally, let's throw in as factor of 100 so that the platform beats ground motion by a factor of 100.  Altogether, the actuator force requirement, which is always of the same order of magnitude as the force requirement, is

(12)(1.5)(6x10-5)(0.88)(100) ~ 10 mN.

Next, the torque requirement.  According to <http://www.iris.edu/hq/instrumentation_meeting/files/pdfs/rotation_iris_igel.pdf>, for a plane shear wave traveling in a medium with phase velocity c, the acceleration a(x, t) is related to the angular rate W'(x, t) through

a(x, t) / W'(x, t) = -2 c.

This implies that |W''(f)| = |a(f)| pi f / c,

where W''(f) is the amplitude spectral density of the angular acceleration and a(f) of the transverse linear acceleration.  I assume that the medium is cement, which according to Wolfram Alpha has a shear modulus of mu = 2.2 GPa and about the density of water: rho ~ 1000 kg/m3.  The shear wave speed in concrete is c = sqrt(mu / rho) ~ 1500 m/s.

The maximum of the acceleration requirement graph is, again, 6x10-5 m/s2 at 500 Hz..  According to Janeen's SolidWorks drawing, the largest principal moment of inertia of the SOS is about 0.26 kg m2.  Including the same fudge factor of (1.5)(100), the net torque requirement is

(0.26) (1.5) (6x10-5) (100) pi (500) / (1500) N m ~ 2.5x10-3 N m.

The quotient of the torque and force requirements is about 0.25 m, so, using some of my previous results, the dimensions of the platform should be as follows:

radius of top plate = 0.25 m,

radius of bottom plate = 2 * 0.25 m = 0.5 m, and

plate separation in home position = sqrt(3) * 0.25 m = 0.43 m.

 

One last thing: perhaps the static load should be taken up directly by the piezos.  If this is the case, then we might rather take the force requirement as being

(10 m/s2)(1.5)(12 kg) = 180 N.

An actuator that can exert a dynamic force of 180 N would easily meet the ground motion requirements by a huge margin.  The dimensions of the platform could also be reduced.  The alternative, I suppose, would be for each piezo to be mechanically in parallel with some sort of passive component to take up some of the static load.

  6194   Thu Jan 12 23:19:56 2012 KojiUpdateSUSc1iscex is fine now

c1iscex is working as before and the optic is damped.


What I checked

1. I went to the X-end rack. I found the io-chassis was turned off.

2. I shutdown c1iscex, turned off, and turned on everything. Again, we did not have any signal from the ADC into c1scx model.

However, I found that c1x01 indicates healthy ADC signals.
This means that the connection between the IOP and the c1scx model was wrong ==> Simulated Plant

3. Burtrestored X'mas eve snapshot. This restored the gains and matrices as well as C1:SCX-SIM_SWITCH
which switches the input between the real ADCs and simulated plant.

4. The signals came back to c1scx.

  6193   Thu Jan 12 23:13:42 2012 KojiConfigurationWIKI-40M UpdateUnable to create Wiki page

Quote:

 I can't create a new page on the 40m wiki.  The page that I was trying to create is

http://blue.ligo-wa.caltech.edu:8000/40m/Stewart

I get this message when I try to save the new page:

Page could not get locked. Unexpected error (errno=13).

This address for wiki is obsolete. Recently it was switched to https://wiki-40m.ligo.caltech.edu/
Jamie is working on automatic redirection from the old wiki to the new place.

The new one uses albert.einstein authentication.

 

ELOG V3.1.3-